SOHO DAILY REPORT Operations Day - 298/2002 (25 October 02) Mission Day 2520 SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC Ref. #: SH-FOT-DR-021025 ==================================================================== PASS #1 Madrid DSS66 0628-1200 PASS #2 Goldstone DSS16 1355-2005 PASS #3 Canberra DSS46 1943-299/0628 ********************************************************************* PASS #1 Madrid DSS66 0628-1200 Anomalies: None NRT: None Activities Summary: - Collected MDI-M (0709-0726) - Collected MDI-H (0726-1140) - SSR SEF Reset - SSR Record ------------------ PASS #2 Goldstone DSS16 1355-2005 Anomalies: G02-0566 1600 Bad clock delta. Deferred troublshooting due to VC2 online and NRT commanding. Performed on next pass. DR#101747. S02-0064 1937 SUMER PLM temperature QTSI1A trips yellow high limit. NRT: 1523 - 1933 Activities Summary: - SSR Dump - Collected MDI-M (1520 - 1933) - Uplinked delayed load VIRA21026, SWA021025101 - Uplinked RSL Table - Reaction Wheel Speed Monitoring Daily Adjustment - CRS Drift Estimation Update - New Guide Star (5.75, slot 4) - New tracking star (6.95 mag, slot 1) - Collect Ranging Data ------------------ PASS #3 Canberra DSS46 1943-299/0620 Anomalies: G02-0567 2006 1 hour ranging lost. Station unable to initiate ranging. Equipment problem. DR#C101799. G02-0568 2055 Wayward commands. First at 2055, 9 retransmissions 7 minute NRT delay. Second at 2114, 6 retransmissions and 5 minute NRT pause. NRT: 2012 - 2245 Activities Summary: - Collected MDI-M (2010-0028) - Collected MDI-H (0028-0600) - OBT Distribution to SWAN, CELIAS - Collected Ranging Data - SSR SEF Reset ------------------ -- End of Report --